OpenStudy (anonymous):

Divide f(x) by d(x), and write a summary statement in the form indicated. f(x) = x4 + 5x3 - 2x2 + 5x - 3; d(x) = x2 + 1

OpenStudy (anonymous):

\[x^4+5 x^3-2 x^2+5 x-3=\left(x^2+1\right) \left(x^2+5 x-3\right) \]Does this help?

OpenStudy (anonymous):

can you explain it to me please?

OpenStudy (anonymous):

The RHS shows the result of factoring f(x) . If both sides are divided by x^2+1, then x^2+5x-2 is the result.

OpenStudy (anonymous):

Synthetic division shows a quotient of x^2 + 5 x - 3 and a remainder of zero.

OpenStudy (anonymous):

thanks a lot